alisha is x years old. jayden is 4 times as old as alisha.
write an algebraic expression for jaydens age.
the sum of both their ages is 45.
write an equation to show this information. solve it.
how old is jayden

Answer by Theo(13342)

alicia is x years old.

jayden is 4 times as old as alisha.

that makes jayden's age equal to 4 * x.

the sum of their ages is 45.

that makes x + 4 * x = 45

combine like terms to get 5 * x = 45.

solve for x to get x = 45/5 = 9.

alisha is 9 years old.

jayden is 4 * 9 = 36 years old.

the sum of their ages if 36 + 9 = 45 years.

solution is confirmed to be good.

the solution is that jayden is 36 years old.